# On the relaxation to equilibrium of a quantum oscillator interacting with a radiation field

### v1

@@ 1 Introduction and outline
In this article we investigate the relaxation to thermodynamical equilibrium of a one-dimensional quantum harmonic oscillator interacting with a radiation field, for instance a quantized electromagnetic field in a vacuum or lattice vibrations in a crystal, acting as a heat bath. More specifically, we are primarily interested in the long time behavior of the occupation probabilities associated with each quantum energy state of the oscillator after its initial interaction with the bath. The system governing the evolution of those probabilities consists of infinitely many master equations, that is, one such equation for each quantum state. The equations we consider are typical gain-loss equations whose structure is already apparent in [18], albeit in a very different context. They represent a Markovian approximation to a more general evolution system that takes the form of an integro-differential equation with a memory term for the reduced density operator under consideration. Accordingly, we organize our article in the following way: In Section 2 we define the class of master equations to be analyzed and prove the self-adjointness of their generator on a dense domain of a suitable Hilbert space. We proceed by proving that the generator in question has a compact resolvent and thereby a discrete spectrum, which requires a concrete realization of its domain as a weighted Sobolev space of sequences compactly embedded in the given Hilbert space. This allows us to prove that the system generates a dynamical system of hyperbolic type, and from there we obtain the spectral resolution of the corresponding evolution semi-group, which is holomorphic and contracting. Knowing this, we then prove that any initial probability distribution remains a probability distribution for all times, which eventually converges exponentially rapidly, as time becomes large, to the Gibbs equilibrium distribution characterized by the temperature of the bath. In Section 3 we determine the spectrum of a linear pencil naturally associated with the system of equations analyzed in Section 2. We conclude the article with an appendix, in which we prove a generalization of the compactness result used in Sections 2 and 3 along with some consequences. Throughout this work we make a point of giving detailed proofs of all statements, with appropriate references whenever necessary.
@@
Needless to say, much is known regarding the properties of a quantum harmonic oscillator in a heat bath and their physical consequences (see, e.g. [5], [12] or [26] and some of the references therein). Thus, the focus of our approach to the problem of relaxation by means of spectral arguments is not so much about the results they imply as it is about the strategy to obtain them. As such, our method may then be considered as complementary to the existing ones.
@@ 2 A spectral theorem for a class of infinitely many Pauli master equations
Regarding the question of how the dynamics of the coupled system consisting of the harmonic oscillator and the radiation field acting as a heat bath leads to (2), we refer the reader for instance to [$12,17,26$-$28$]. In a nutshell, as in [12] and in Sections 3 and 4 of Chapter XVII in [26], the bath is modelled by an infinite system of quantum harmonic oscillators whose frequencies satisfy certain conditions. The time evolution of the density operator describing that coupling is then governed by the laws of Quantum Mechanics. Subsequently, various methods are used to eliminate the bath variables to eventually obtain the evolution equation of the so-called reduced density operator, that is, the density operator of the given oscillator after its initial interaction with the bath. Among others, one of those methods is the celebrated Projection Technique devised in [$17,27,28$]. As a consequence and within a Markovian approximation, equation (2) emerges naturally with the functions $t\rightarrow$ $p_{\mathsf{m}}(t)$ being defined as the diagonal matrix elements of the reduced density operator with respect to the $\mathsf{m}^{\mathsf{th}}$ Hermite function. We refer the reader for instance to [$5$-$8$] for a rigorous implementation of some of those formal developments.
@@
$\left\|R\mathsf{f}_{\mathsf{m}}\right\|_{2,w_{\beta}}^{2}$ $=$ $\rho^{2}\left\{(\mathsf{m}+\exp\left[-\beta\right](\mathsf{m}+1)^{2}+\exp\left[-\beta\right]\left(2\mathsf{m}^{2}+2\mathsf{m}+1\right)\right\}$
@@
Thus, it remains to examine the consequences of the recurrence relation (30). We first consider the case where $\mathsf{q}\in D(R^{\ast})$ is real. With our choice of $\varepsilon$ this leads to
$q_{\mathsf{m}}=\exp\left[-\mathsf{m}\beta\right]q_{\mathsf{0}}$
for every $\mathsf{m}\in\mathbb{N}$, which gives (24) as in Proposition 3 but with some $q_{0}\in\mathbb{R}$. The general case for $\mathsf{q=r+}i\mathsf{s}$ with $\mathsf{r,s}\in D(R^{\ast})$ both real then follows from complexification. Indeed we have
@@ 3 On the spectral properties of a linear pencil associated with $R^{\ast}$
We conclude this article with the following:
Appendix. On the compactness of the embedding $h_{\mathbb{C},w_{\beta}}^{1}\hookrightarrow l_{\mathbb{C},w_{\beta}}^{2}$
@@
Acknowledgments: The author would like to thank the Fundação para a Ciência e a Tecnologia of the Português Government (FCT) for its financial support under grant UIDB/04561/2020.

### v2

@@ 1 Introduction and outline
In this article we investigate the relaxation to thermodynamical equilibrium of a one-dimensional quantum harmonic oscillator interacting with a radiation field, for instance a quantized electromagnetic field in a vacuum or lattice vibrations in a crystal, acting as a heat bath. More specifically, we are primarily interested in the long time behavior of the occupation probabilities associated with each quantum energy state of the oscillator after its initial interaction with the bath. The system governing the evolution of those probabilities consists of infinitely many master equations, that is, one such equation for each quantum state. The equations we consider are typical gain-loss equations whose structure is already apparent in [19], albeit in a very different context. They represent a Markovian approximation to a more general evolution system that takes the form of an integro-differential equation with a memory term for the reduced density operator under consideration. Accordingly, we organize our article in the following way: In Section 2 we define the class of master equations to be analyzed and prove the self-adjointness of their generator on a dense domain of a suitable Hilbert space. We proceed by proving that the generator in question has a compact resolvent and thereby a discrete spectrum, which requires a concrete realization of its domain as a weighted Sobolev space of sequences compactly embedded in the given Hilbert space. This allows us to prove that the system generates a dynamical system of hyperbolic type, and from there we obtain the spectral resolution of the corresponding evolution semi-group, which is holomorphic and contracting. Knowing this, we then prove that any initial probability distribution remains a probability distribution for all times, which eventually converges exponentially rapidly, as time becomes large, to the Gibbs equilibrium distribution characterized by the temperature of the bath. In Section 3 we determine the spectrum of a linear pencil naturally associated with the system of equations analyzed in Section 2. We conclude the article with two appendices. In the first one we prove a generalization of the compactness result used in Sections 2 and 3 along with some consequences. In the second one we briefly explain how the interaction of the harmonic oscillator with the radiation field leads to the considered system of master equations, starting out with the evolution equation for the reduced density operator of the harmonic oscillator after its initial interaction with the heat bath and the elimination of the bath variables.
Throughout this work we make a point of giving detailed proofs of all statements, with appropriate references whenever necessary.
@@
Needless to say, much is known regarding the properties of a quantum harmonic oscillator in a heat bath and their physical consequences (see, e.g. [5], [12] or [27] and some of the references therein). Thus, the focus of our approach to the problem of relaxation by means of spectral arguments is not so much about the results they imply as it is about the strategy to obtain them. Indeed, to the best of our knowledge, the spectral analysis of the generator of the master equations viewed as an unbounded self-adjoint operator on some suitable weighted sequence space and based on compactness arguments is new, as is its relation to the spectral theory of linear pencils. As such, our method may then be considered as complementary to the existing ones, particularly to those set forth in [5] and [6], which make use of completely different procedures, and thus renders our contribution self-contained.
@@ 2 A spectral theorem for a class of infinitely many Pauli master equations
Regarding the question of how the dynamics of the coupled system consisting of the harmonic oscillator and the radiation field acting as a heat bath leads to (2), we refer the reader for instance to [$12,17,26$-$28$], and to Appendix B for a brief summary. In a nutshell, as in [12] and in Sections 3 and 4 of Chapter XVII in [27], the bath is modelled by an infinite system of quantum harmonic oscillators whose frequencies satisfy certain conditions. The time evolution of the density operator describing that coupling is then governed by the laws of Quantum Mechanics. Subsequently, various methods are used to eliminate the bath variables to eventually obtain the evolution equation of the so-called reduced density operator, that is, the density operator of the given oscillator after its initial interaction with the bath. Among others, one of those methods is the celebrated Projection Technique devised in [$17,27,28$]. As a consequence and within a Markovian approximation, equation (2) emerges naturally with the functions $t\rightarrow$ $p_{\mathsf{m}}(t)$ being defined as the diagonal matrix elements of the reduced density operator with respect to the $\mathsf{m}^{\mathsf{th}}$ eigenfunction of the oscillator. We refer the reader for instance to [$5$-$8$] for a rigorous implementation of some of those formal developments.
@@
$\left\|R\mathsf{f}_{\mathsf{m}}\right\|_{2,w_{\beta}}^{2}$ $=$ $\rho^{2}\left\{(\mathsf{m}+\exp\left[-\beta\right](\mathsf{m}+1))^{2}+\exp\left[-\beta\right]\left(2\mathsf{m}^{2}+2\mathsf{m}+1\right)\right\}$
@@
Thus, it remains to examine the consequences of the recurrence relation (30), which implies
$\left|q_{\mathsf{m}}\right|=\exp\left[-\beta\mathsf{m}\right]\left|q_{\mathsf{0}}\right|$
for every $\mathsf{m}\in\mathbb{N}$ with our choice of $\varepsilon$. If $\mathsf{q}\in D(R^{\ast})$ is real, the preceding expression gives
$q_{\mathsf{m}}=\exp\left[-\beta\mathsf{m}\right]\left(\pm q_{\mathsf{0}}\right):=\exp\left[-\beta\mathsf{m}\right]\hat{q}_{\mathsf{0}},$
hence (24) as in Proposition 3 but with some parameter $\hat{q}_{0}\in\mathbb{R}$. The general case for $\mathsf{q=r+}i\mathsf{s}$ with $\mathsf{r,s}\in D(R^{\ast})$ both real then follows from complexification. Indeed we have
@@ 3 On the spectral properties of a linear pencil associated with $R^{\ast}$
We conclude this article with two appendices:
Appendix A. On the compactness of the embedding $h_{\mathbb{C},w_{\beta}}^{1}\hookrightarrow l_{\mathbb{C},w_{\beta}}^{2}$
@@
Appendix B. On the relation between quantum dynamics and Pauli master equations for the harmonic oscillator
In this last appendix we briefly and informally sketch the way in which the dynamics of a quantum harmonic oscillator out of equilibrium in contact with a radiation field leads to master equations of the form (1) or (4). In so doing we refer the reader to the literature listed in the main part of the text, more specifically to Sections 2, 3 and 4 in Chapter XVII of [27]. Radiation fields such as quantized electrodynamic fields in a vacuum or lattice vibrations in a crystal are typically modelled by an infinite system of quantum harmonic oscillators. According to the laws of Quantum Mechanics, the equation governing the time evolution of the density operator $\mathcal{R}$ of a single harmonic oscillator in contact with such a field is then of the form
$\frac{d\mathcal{R}\left(t\right)}{dt}=-i\left[H_{O}+H_{B}+\gamma H_{I},\mathcal{R}\left(t\right)\right]_{-}$
in suitable units. In the preceding expression, $H_{O}$ stands for the Hamiltonian operator of the single harmonic oscillator, $H_{B}$ for that of the field acting as a bath while $H_{I}$ denotes the interaction Hamiltonian between the two, $\gamma\in\mathbb{R}$ being a coupling constant. All four operators are related by the usual commutator $\left[.\right]_{-}$. More specifically,
$H_{O}=a_{0}^{\ast}a_{0},$ (89)
$H_{B}=\sum_{\mathsf{n}\neq 0}\omega_{\mathsf{n}}a_{\mathsf{n}}^{\ast}a_{\mathsf{n}}$ (90)
and
$H_{I}=\sum_{\mathsf{n}\neq 0}\gamma_{\mathsf{n}}\left(a_{\mathsf{n}}a_{0}^{\ast}+a_{\mathsf{n}}^{\ast}a_{0}\right)$
in suitable units, where $a_{0},a_{0}^{\ast}$ and $a_{\mathsf{n}},a_{\mathsf{n}}^{\ast}$ are the usual annihilation and creation operators for $H_{O}$ and $H_{B}$, respectively, $\omega_{\mathsf{n}}$ the frequencies of the bath oscillators and $\gamma_{\mathsf{n}}$ some additional coupling constants. In (89) and (90) the ground state energy is omitted as it is irrelevant for the time evolution, while $H_{O}$ has been renormalized in such a way that its spectrum corresponds to the choice made in Section 2 immediately after the proof of Lemma 1, that is to say the set of non-negative integers. Assuming that the bath is in thermodynamical equilibrium at inverse temperature $\beta>0$ and putting the oscillator in contact with it at the initial time $t=0$, it is then possible by means of several simplifying hypotheses to eliminate the bath variables. Among others, those hypotheses include the weak coupling limit and the Markovian approximation, in which case one eventually obtains in closed form the so-called quantum master equation for the reduced density operator $R_{O}$ of the oscillator, namely
$\frac{d\mathcal{R}_{O}\left(t\right)}{dt}$
$=$ $-i\left[a_{0}^{\ast}a_{0},\mathcal{R}_{O}\left(t\right)\right]_{-}$
$+\sigma\left(a_{0}^{\ast}\mathcal{R}_{O}\left(t\right)a_{0}-\frac{1}{2}a_{0}a_{0}^{\ast}\mathcal{R}_{O}\left(t\right)-\frac{1}{2}\mathcal{R}_{O}\left(t\right)a_{0}a_{0}^{\ast}\right)$
$+\rho\left(a_{0}\mathcal{R}_{O}\left(t\right)a_{0}^{\ast}-\frac{1}{2}a_{0}^{\ast}a_{0}\mathcal{R}_{O}\left(t\right)-\frac{1}{2}\mathcal{R}_{O}\left(t\right)a_{0}^{\ast}a_{0}\right).$
In (3), the constants $\sigma$ and $\rho$ are such that
$\frac{\sigma}{\rho}=\frac{1-\exp\left[-\beta\right]}{\exp\left[\beta\right]-1}=\exp\left[-\beta\right],$
which turn out to be exactly those that appear in (3) and (9). In fact, it is now easy to derive (4) from (3). Let $\left|\mathsf{m}\right\rangle$ be the eigenstate of the harmonic oscillator correponding to its $\mathsf{m}^{\mathsf{th}}$ energy level, with $\mathsf{m}\in\mathbb{N}$, and let us recall that
$a_{0}\left|\mathsf{m}\right\rangle$ $=$ $\sqrt{\mathsf{m}}\left|\mathsf{m-1}\right\rangle,$
$a_{0}^{\ast}\left|\mathsf{m}\right\rangle$ $=$ $\sqrt{\mathsf{m+1}}\left|\mathsf{m+1}\right\rangle,$ (92)
$a_{0}^{\ast}a_{0}\left|\mathsf{m}\right\rangle$ $=$ $\mathsf{m}\left|\mathsf{m}\right\rangle,$
and that $a_{0}^{\ast}$ is the formal adjoint of $a_{0}$ (see, e.g., Chapter XII in [16] for a very clear derivation of these facts). Defining the probability of occupation of the $\mathsf{m}^{\mathsf{th}}$ level at time $t$ as
$p_{\mathsf{m}}(t)=\left\langle\mathsf{m}\right|\mathcal{R}_{O}\left(t\right)\left|\mathsf{m}\right\rangle,$
which indeed satisfies $0\leq p_{\mathsf{m}}(t)\leq 1$ by virtue of the standard properties of density operators, we then get successively from (3) and repeated applications of all three equations in (92) the relation
$\frac{dp_{\mathsf{m}}(t)}{dt}$
$=$ $-i\left\langle\mathsf{m}\right|\left[a_{0}^{\ast}a_{0},\mathcal{R}_{O}\left(t\right)\right]_{-}\left|\mathsf{m}\right\rangle$
$+\sigma\left(\left\langle\mathsf{m}\right|a_{0}^{\ast}\mathcal{R}_{O}\left(t\right)a_{0}\left|\mathsf{m}\right\rangle-\frac{1}{2}\left\langle\mathsf{m}\right|a_{0}a_{0}^{\ast}\mathcal{R}_{O}\left(t\right)\left|\mathsf{m}\right\rangle-\frac{1}{2}\left\langle\mathsf{m}\right|\mathcal{R}_{O}\left(t\right)a_{0}a_{0}^{\ast}\left|\mathsf{m}\right\rangle\right)$
$+\rho\left(\left\langle\mathsf{m}\right|a_{0}\mathcal{R}_{O}\left(t\right)a_{0}^{\ast}\left|\mathsf{m}\right\rangle-\frac{1}{2}\left\langle\mathsf{m}\right|a_{0}^{\ast}a_{0}\mathcal{R}_{O}\left(t\right)\left|\mathsf{m}\right\rangle-\frac{1}{2}\left\langle\mathsf{m}\right|\mathcal{R}_{O}\left(t\right)a_{0}^{\ast}a_{0}\left|\mathsf{m}\right\rangle\right)$
$=$ $\sigma\left(\mathsf{m}p_{\mathsf{m-1}}(t)-\left(\mathsf{m+1}\right)p_{\mathsf{m}}(t)\right)+\rho\left(\left(\mathsf{m+1}\right)p_{\mathsf{m+1}}(t)-\mathsf{m}p_{\mathsf{m}}(t)\right).$
But regrouping terms differently in the last line we see that this is exactly (4) or, equivalently, the system of master equations (1) with the rates (3).
@@ Appendix B. On the relation between quantum dynamics and Pauli master equations for the harmonic oscillator
Acknowledgments: The author would like to thank the referee for his or her suggestions which helped improve the presentation of this article. He also thanks the Fundação para a Ciência e a Tecnologia of the Português Government (FCT) for its financial support under grant UIDB/04561/2020.